OpenStudy (anonymous):

There are three tunnels A, B, and C, and the length of the tunnels are in the proportion of 2:4:1, respectively. If a train takes 16 seconds to travel through tunnel A and 30 seconds to travel through tunnel B, how long does it take to travel through tunnel C?

OpenStudy (anonymous):

any ideas guys?

OpenStudy (anonymous):

its a ratio question.

OpenStudy (anonymous):

u know how to do it?

OpenStudy (anonymous):

think of the ratio as lengths

OpenStudy (anonymous):

2:4:1

OpenStudy (anonymous):

length A=? B=? C=?...

OpenStudy (anonymous):

like A =2, B=4, C=1?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

but I don't get the time part?

OpenStudy (anonymous):

with the seconds

OpenStudy (anonymous):

v=d/t so if velocity is the same lets say v=1. solve for d

OpenStudy (anonymous):

because we want to find the distance of tunnel c

OpenStudy (anonymous):

how do u solve for d? what's the value for t?

OpenStudy (anonymous):

i got of track a bit we want to find the time to travel c and if v*d=t and v=1 t=?

OpenStudy (anonymous):

so if you know the ratio of the distances you also know the ratio of the ....?

OpenStudy (anonymous):

the time?

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

replace A time= B time = C time = and what the ratio between them.

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

how does that work though? the time of A: B is 16:30, which is not 2:4

OpenStudy (anonymous):

or double the time i assume it is using significant figures. so ~2x16=30 or its a book typo because a train that slows down and speeds up would make this problem very hard.

OpenStudy (anonymous):

so the answer should be what?

OpenStudy (anonymous):

are you sure its 16 not 15?

OpenStudy (anonymous):

or 32 not 30.?

OpenStudy (anonymous):

it's on my paper

OpenStudy (anonymous):

anyway, the answer would be either 15 or 16 right?

OpenStudy (anonymous):

1/4 of 32 =

OpenStudy (anonymous):

ok sorry we assumed that v = 1

OpenStudy (anonymous):

so we have to solve for v

OpenStudy (anonymous):

not equal to 1

OpenStudy (anonymous):

oops I meant 7.5 or 8 right?

OpenStudy (anonymous):

right but i think 30 is wrong and should be 32 so its 8

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can u help me with some more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Dad is now 43 years old and son is now 13. How many years will pass so that dad's age is twice his son's age?

OpenStudy (anonymous):

I know it's 17 but how do I do it the algebraic way?

OpenStudy (anonymous):

so

OpenStudy (anonymous):

we have two variables D=dad's age and S= sons age

OpenStudy (anonymous):

and a ratio

OpenStudy (anonymous):

D= 2S right?

OpenStudy (anonymous):

does this make sense

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

there is one unknown n = number of years that between know and when the ratio holds.

OpenStudy (anonymous):

and they each age by n

OpenStudy (anonymous):

(43+n)=2(13+n) solve for n

OpenStudy (anonymous):

(D+nyears)/(S +nyears)=2

OpenStudy (anonymous):

ur so smart :)

OpenStudy (anonymous):

haha cheers. I got to go have a good day.

OpenStudy (anonymous):

2 more questions, is that ok?

OpenStudy (anonymous):

it will be quick

OpenStudy (anonymous):

I cant man sorry. good luck.